Air bags in cars inflate when an elect
rical spark activates sodium azide (NaN₃) so that it decomposes to sodium metal and nitrogen gas. How many liters of N₂ gas are formed for each gram of NaN₃ consumed, if the reaction is carried out at STP.
Answer #1

Answer:-

Firstly a balanced decomposition reaction is written then using the simple concept of stoichiometry, moles of nitrogen are calculated then volume is calculated at STP.

The answer is given in the image,

Answer: 2 MaM3 — 212 + 3 M30g) Molaimana Nadz = 23+ 3814 = 65 gomol. So 65 g Nam3 = 3 mol Na . 1 g Nana = 3 melna = 0.04 6154